Terry rolls a number cube labeled 1 through 6 two times . What is the probability that the number cube will land on 5 for the first roll and 3 for the second roll ?

Answer by stanbon(75887)

Terry rolls a number cube labeled 1 through 6 two times . What is the probability that the number cube will land on 5 for the first roll and 3 for the second roll ?
-------
The events are independent.
P(5 and 3) = P(5)*P(3) = 1/6 * 1/6 = 1/36
=============
Cheers,
Stan H.